# Impact of general relativistic accretion on primordial black holes

### v1

@@ I Introduction
The groundbreaking discovery of gravitational waves by the LIGO and Virgo collaborations in 2016 LIGOScientific:2016aoc further propelled BHs into the spotlight, intensifying their relevance in observational cosmology. In this realm, primordial black holes (PBHs) stand out as a potential key to unlocking several cosmological puzzles. These PBHs, formed from the gravitational collapse of overdensities in the early universe, span a wide mass range — from sub-gram scales to solar masses Carr & Hawking (1974); Carr (1975); Carr:2020gox ; Auffinger:2022khh ; Khlopov:2008qy . Depending on their mass, PBHs can undergo both accretion and evaporation during cosmic evolution, influencing various phenomena such as dark matter (DM) production Masina:2020xhk ; Hooper:2019gtx ; Baker:2022rkn ; Cheek:2021odj ; Calabrese:2021src ; Chaudhuri:2023aiv , baryogenesis Baumann:2007yr ; Fujita:2014hha ; Datta:2020bht ; Hook:2014mla ; Barman:2022gjo , and cosmic microwave background (CMB) distortions Tashiro:2008sf ; Acharya:2020jbv ; Ricotti:2007au ; Ali-Haimoud:2016mbv ; Piga:2022ysp ; Poulin:2017bwe ; Serpico:2020ehh ; Ziparo:2022fnc to name a few. Accretion is one of the pivotal processes of PBHs evolution in the early universe Michel (1972); Harada:2004pf ; Harada:2004pe ; Lora-Clavijo:2013aya ; Mack:2006gz ; DeLuca:2020fpg ; DeLuca:2021pls ; DeLuca:2023bcr ; Yang:2022puh ; Yuan:2023bvh ; Zhang:2023rnp . While early studies on PBH accretion primarily focused on Bondi-Hoyle Ricotti:2007jk ; Tabasi:2021cxo ; Tabasi:2022fap ; Stojkovic:2004hz ; Jangra:2024sif ; Nayak:2009wk ; Park:2010yh ; Wang:2023qxj ; Manzoni:2023gon ; khlopov:hal-03129467 models in the Newtonian framework, the general relativistic effect could be particularly important for PBHs that are embedded in a hot and dense environment of relativistic plasma in the early universe.
Motivated by this, we present a reformulation of the PBH accretion process within a general relativistic framework—a regime that, to the best of our knowledge, remains largely unexplored, and investigate its impact on the evolution of PBHs for the first time. Our analysis shows that relativistic corrections to the Bondi accretion model significantly enhance the accretion rate, leading to substantial mass growth of ultralight PBHs prior to their evaporation via Hawking radiation. By numerically solving the coupled equations governing accretion and evaporation in an expanding Universe, we derive updated constraints on the initial PBH mass spectrum, refine bounds on their abundance, and assess their viability as dark matter candidates. In addition, we predict a measurable shift in the high-frequency gravitational wave spectrum due to accretion-induced mass evolution. This unified treatment of PBH dynamics incorporating both evaporation and relativistic accretion provides a more accurate and testable framework for PBH based dark matter scenarios in light of current cosmological data.
@@ II General relativistic accretion
The horizon is defined as $g^{rr}=1/g_{rr}=0$ and the metric coefficients are functions of $(r,\theta)$. In this background, the dynamics of an ideal fluid is described by the conservation equations, $\nabla_{\nu}T^{\mu\nu}=0~{};~{}\nabla_{\mu}j^{\mu}=0$, where the indices $\mu$ and $\nu$ run from 00 to $3$. The symbol $\nabla_{\mu}$ denotes the covariant derivative. Recall that any ideal fluid of energy density $\rho$, pressure $p$, and mass density $\tilde{\rho}$ is characterized by the energy momentum tensor $T^{\mu\nu}$, and the particle number current $j^{\mu}$, which are expressed as,
@@
In equation (3), the roman index assumes $i=1,2,3$. Conventionally, the evaluation of the accreting matter is described in terms of the quantity defined in co-rotating frame of the fluid considering the following series of variable transformations: Azimuthal velocity is defined as $v_{\phi}^{2}=(u^{\phi}u_{\phi})/(-u^{t}u_{t})$ with the associated Lorentz factor $\gamma_{\phi}=1/\sqrt{(1-v^{2}_{\phi})}$; the polar velocity is defined as $v_{\theta}^{2}=\gamma^{2}_{\phi}(u^{\theta}u_{\theta})/(-u^{t}u_{t})$ with the associated Lorentz factor $\gamma_{\theta}=1/\sqrt{(1-v^{2}_{\theta})}$, and finally the radial velocity is defined as $v^{2}=\gamma^{2}_{\phi}\gamma^{2}_{\theta}(u^{r}u_{r})/(-u^{t}u_{t})$ with the associated radial Lorentz factor $\gamma_{v}=1/\sqrt{(1-v^{2})}$. All detailed equations formulated using the co-rotating fluid variables are provided in the reference Dihingia:2018tlr . However, for our present study we focus on spherical accretion and assume $v_{\theta}=v_{\phi}=0$. Expanding the Eq. (3) for $i=r$, we get,
@@
Here, $\Phi^{\text{eff}}~{}[=\frac{1}{2}\ln(-g_{tt})]$ is the general relativistic effective potential and $h~{}[=\rho+p)/\tilde{\rho}]$ denotes the enthalpy of the flow. In this work, we consider Schwarzschild BH where $g_{tt}=-g^{rr}=-f(r),~{}g_{t\phi}=0,~{}g_{\theta\theta}=g_{\phi\phi}/\sin^{2}\theta=r^{2}$ with $f(r)=1-{2GM}/{r}$. Here, $M$ denotes the mass of the BH. The effective gravitational potential perceived by the fluid is given by $\Phi^{\text{eff}}=\frac{1}{2}\ln f(r)$, and the radial momentum equation transforms into,
@@
Equations (6) and (7) are the two master equations that govern the spherical accretion dynamics for an asymptotically flat BH. However, to understand accretion onto PBHs, we reformulate these equations in the Friedmann–Lemaître–Robertson–Walker background, parametrized by the scale factor $a(t)$, with the metric given by $ds^{2}=-dt^{2}+a^{2}(t)(dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta\,d\phi^{2})$. For this, we follow the conventional approach detailed in Ricotti:2007jk .
@@ II.1 Accretion in the expanding background
where $H~{}(=\dot{a}/a)$ is the Hubble parameter and $v_{p}~{}(=a\dot{x})$. Incorporating Eq. (25) into the accretion equations, and integrating Eq. (6) we get,
@@
The integration constant is set by the condition that far away from the BH ($x\rightarrow\infty$), $v\rightarrow 0$, and $\rho\rightarrow\rho_{\infty}\propto a^{-3(1+\omega)}$, which corresponds to the background energy density of the Universe, characterized by the generic isothermal equation of state $p=\omega\rho$. The Hubble parameter satisfies the dynamical equation $H^{2}=\dot{a^{2}}/a^{2}=({1}/{3M_{p}^{2}})\rho$, which yields $H=H_{\rm in}\left(a/a_{\rm in}\right)^{-\frac{3}{2}(1+\omega)}$, with index “${\rm in}$” being some initial time which we identify later. Here, $M_{p}=1/\sqrt{8\pi G}\simeq 2.435\times 10^{18}\text{Gev}$, being the reduced Planck mass. Utilizing this, the relation between energy density and mass density of the fluid can be obtained by integrating the enthalpy generation Eq. (4) as,
@@
In Fig. 1, we have depicted the various representative solution topologies for the inflowing matter. However, the physical solutions are the ones where the inflowing matters pass through a critical point and transits smoothly from subsonic to supersonic branch. At the critical point, one needs to satisfy $N=D=0$ condition simultaneously. Solving this, we get the critical point location as $x_{c}=GM(1+3\omega)/(2a\omega)$ and matter velocity at $x_{c}$ as $v^{c}_{p}=\sqrt{\omega}-Hax_{c}$. Inserting these in Eq. (11), the accretion rate at the critical point becomes,
@@
where $\lambda_{c}=(1/4)\omega^{-3/2}(1+3\omega)^{(1+3\omega)/2\omega}$. Note that in the non-relativistic limit, the critical value reduces to $\lambda_{c}^{\rm NR}=(1/4)\text{exp}(3/2)\omega^{-3/2}$ Ali-Haimoud:2016mbv . Integrating Eq. (13), we obtain
$\frac{M}{M_{\mathrm{in}}}$ $=$ $\left[1-\frac{\lambda_{c}\gamma}{2(1+\omega)}\left(\frac{\frac{3}{2}H_{\rm in}(1+\omega)(t-t_{\rm in})}{\frac{3}{2}H_{\rm in}(1+\omega)(t-t_{\rm in})+1}\right)\right]^{-1},$ (14)
where $M_{\mathrm{in}}~{}(=4\pi\gamma M_{p}^{2}H_{\rm in}^{-1})$ is the initial mass of a PBH with $H_{\rm in}$ being the Hubble parameter at the formation time $t=t_{\rm in}$. Here, $\gamma~{}(=\omega^{3/2})$ is the collapse fraction to form PBH Carr & Hawking (1974). From Eq. (14), we indeed see that PBH gains mass due to accretion and quickly saturates into (see left panel of Fig. 2),
@@ III Hawking evaporation
For example, assuming a PBH with a mass of 10 g formed in a radiation-dominated background ($\omega=1/3$), it completely evaporates at approximately $\tau_{\rm BH}^{\rm ev}\simeq 8.8\times 10^{17}M_{p}^{-1}$. This highlights the wide separation between the timescales of the two different physical processes of our interest, namely accretion and evaporation ($\tau_{\rm BH}^{\rm ev}\gg\tau_{\rm BH}^{\rm acc}$), as depicted in the left panel of Fig. 2.
@@
$N_{j}\simeq\begin{cases}10^{8}\left(\frac{M_{\rm acc}}{1\;\rm g}\right)^{2}~{}~{}~{}~{}~{}\text{for}~{}m_{j}\lesssim T_{\rm BH}^{\rm acc},\\ 10^{14}\left(\frac{10^{10}\text{Gev}}{m_{j}}\right)^{2}~{}\text{for}~{}m_{j}\gtrsim T_{\rm BH}^{\rm acc},\\ \end{cases}$ (20)
@@ IV PBH mass evolution
The evolution of PBH mass, accounting for both relativistic (blue) and non-relativistic (red) accretion in conjunction with Hawking evaporation, is illustrated in the left panel of Fig. 2. Results are presented for two initial masses, namely $10\,\mathrm{g}$ (dashed lines) and $10^{8}\,\mathrm{g}$ (solid lines), respectively.
@@ V BBN Constraint
PBHs evaporating during the era of Big Bang Nucleosynthesis (BBN) can inject additional relativistic degrees of freedom, potentially altering the primordial element abundances predicted by BBN. To preserve the observed consistency of BBN results, PBHs must evaporate before the onset of BBN. This requirement immediately suggests that the PBH mass within the range of $10^{8}\,\text{g}\lesssim M_{\rm BH}\lesssim 10^{13}\,\text{g}$ Keith:2020jww should not form in the early universe. However, when PBH accretion is taken into account, these mass constraints are naturally altered.
@@
$T_{\rm ev}\simeq\sqrt{\frac{3\epsilon}{2}}\left(\frac{\pi^{2}g_{\ast}}{30}\right)^{-1/4}\left(\frac{M_{p}}{M_{\mathrm{in}}}\right)^{3/2}\left(1-\frac{3\gamma\lambda_{c}}{8}\right)^{3/2}M_{p},$ (25)
@@
In the above expression, we consider $g_{*}=106.75$ at the time of PBH formation. Equation (26) clearly demonstrates the impact of accretion on the BBN limit for the minimum PBH mass. For example, if we set $\lambda_{c}=0$ (without accretion), we recover the well known BBN bound on PBH initial mass $M_{\rm in}\simeq 3.94\times 10^{8}\,\mathrm{g}$. On contrary, relativistic accretion ($\lambda_{c}\simeq 10.4$) clearly shifts such bound to one order lower value, $M_{\rm in}\simeq 8.6\times 10^{7}\,\mathrm{g}$. Our analytic estimates match well with the numerical ones illustrated in the right panel of Fig. 2, and further depicts the dependence of the initial PBH mass on $T_{\rm ev}$ for different accretion scenarios.
Figure 2: Left panel: PBH mass ($M$) evolution as a function of scale factor ($a$) for initial masses of $10$ g (dashed) and $10^{8}$ g (solid), comparing relativistic (green) and non-relativistic (red) accretion along with Hawking evaporation. Right panel: Initial PBH mass ($M_{\rm in}$) as a function of evaporation temperature ($T_{\rm ev}$), shown for three cases namely, blue, red, and green curves representing without, with non-relativistic, and with relativistic accretion respectively.

### v2

@@ I Introduction
The groundbreaking discovery of gravitational waves by the LIGO and Virgo collaborations in 2016 LIGOScientific:2016aoc further propelled BHs into the spotlight, intensifying their relevance in observational cosmology 2007coaw.book…..C . In this realm, primordial black holes (PBHs) stand out as a potential key to unlocking several cosmological puzzles. These PBHs, formed from the gravitational collapse of overdensities in the early universe, span a wide mass range — from sub-gram scales to solar masses Carr & Hawking (1974); Carr (1975); Carr:2020gox ; Auffinger:2022khh ; Khlopov:2008qy . Depending on their mass, PBHs can undergo both accretion and evaporation during cosmic evolution, influencing various phenomena such as dark matter (DM) production Masina:2020xhk ; Hooper:2019gtx ; Baker:2022rkn ; Cheek:2021odj ; Calabrese:2021src ; Chaudhuri:2023aiv , baryogenesis Baumann:2007yr ; Fujita:2014hha ; Datta:2020bht ; Hook:2014mla ; Barman:2022gjo ; Hamada:2016jnq , and cosmic microwave background (CMB) distortions Tashiro:2008sf ; Acharya:2020jbv ; Ricotti:2007au ; Ali-Haimoud:2016mbv ; Piga:2022ysp ; Poulin:2017bwe ; Serpico:2020ehh ; Ziparo:2022fnc to name a few. Accretion is one of the pivotal processes of PBHs evolution in the early universe Michel (1972); Harada:2004pf ; Harada:2004pe ; Lora-Clavijo:2013aya ; Mack:2006gz ; DeLuca:2020fpg ; DeLuca:2021pls ; DeLuca:2023bcr ; Yang:2022puh ; Yuan:2023bvh ; Zhang:2023rnp . While early studies on PBH accretion primarily focused on Bondi-Hoyle Ricotti:2007jk ; Tabasi:2021cxo ; Tabasi:2022fap ; Stojkovic:2004hz ; Jangra:2024sif ; Nayak:2009wk ; Park:2010yh ; Wang:2023qxj ; Manzoni:2023gon ; khlopov:hal-03129467 models in the Newtonian framework, the general relativistic effect could be particularly important for PBHs that are embedded in a hot and dense environment of relativistic plasma in the early universe.
Motivated by this, we present a reformulation of the PBH accretion process within a general relativistic framework, and we systematically investigate its impact on the cosmological evolution of PBHs. Our analysis shows that relativistic corrections to the Bondi accretion model significantly enhance the accretion rate, leading to substantial mass growth of ultralight PBHs prior to their evaporation via Hawking radiation. By numerically solving the coupled equations governing accretion and evaporation in an expanding Universe, we derive updated constraints on the initial PBH mass spectrum, refine bounds on their abundance, and assess their viability as dark matter candidates. In addition, we predict a measurable shift in the high-frequency gravitational wave spectrum due to accretion-induced mass evolution. This unified treatment of PBH dynamics incorporating both evaporation and relativistic accretion provides a more accurate and testable framework for PBH based dark matter scenarios in light of current cosmological data.
@@ II General relativistic accretion
The horizon is defined as $g^{rr}=1/g_{rr}=0$ and the metric coefficients are functions of $(r,\theta)$. In this background, the dynamics of an ideal fluid is described by the conservation equations, $\nabla_{\nu}T^{\mu\nu}=0~;~\nabla_{\mu}j^{\mu}=0$, where the indices $\mu$ and $\nu$ run from 0 to $3$. The symbol $\nabla_{\mu}$ denotes the covariant derivative. Recall that any ideal fluid of energy density $\rho$, pressure $p$, and mass density $\tilde{\rho}$ is characterized by the energy momentum tensor $T^{\mu\nu}$, and the particle number current $j^{\mu}$, which are expressed as,
@@
In equation (3), the roman index assumes $i=1,2,3$. Conventionally, the evaluation of the accreting matter is described in terms of the quantity defined in co-rotating frame of the fluid considering the following series of variable transformations: Azimuthal velocity is defined as $v_{\phi}^{2}=(u^{\phi}u_{\phi})/(-u^{t}u_{t})$ with the associated Lorentz factor $\gamma_{\phi}=1/\sqrt{(1-v^{2}_{\phi})}$; the polar velocity is defined as $v_{\theta}^{2}=\gamma^{2}_{\phi}(u^{\theta}u_{\theta})/(-u^{t}u_{t})$ with the associated Lorentz factor $\gamma_{\theta}=1/\sqrt{(1-v^{2}_{\theta})}$, and finally the radial velocity is defined as $v^{2}=\gamma^{2}_{\phi}\gamma^{2}_{\theta}(u^{r}u_{r})/(-u^{t}u_{t})$ with the associated radial Lorentz factor $\gamma_{v}=1/\sqrt{(1-v^{2})}$. All detailed equations formulated using the co-rotating fluid variables are provided in the reference Dihingia:2018tlr . However, for our present study we focus on spherical accretion and assume $v_{\theta}=v_{\phi}=0$ 1984ApJ…277..296H . Expanding the Eq. (3) for $i=r$, we get,
@@
Here, $\Phi^{\text{eff}}~[=\frac{1}{2}\ln(-g_{tt})]$ is the general relativistic effective potential and $h~[=\rho+p)/\tilde{\rho}$] denotes the enthalpy of the flow. In this work, we consider Schwarzschild BH where $g_{tt}=-g^{rr}=-f(r),~g_{t\phi}=0,~g_{\theta\theta}=g_{\phi\phi}/\sin^{2}\theta=r^{2}$ with f(r)=12GM/rf(r)=1-{2GM}/{r}. Here, $M$ denotes the mass of the BH. The effective gravitational potential perceived by the fluid is given by $\Phi^{\text{eff}}=\frac{1}{2}\ln f(r)$, and the radial momentum equation transforms into,
@@
Together, Eqs. (6) and (7) constitute the fundamental set of equations that describe the dynamics of steady-state, spherical accretion onto a BH in an asymptotically flat spacetime.
In the non-relativistic (Newtonian) limit, valid at large radial distances from the source, the fluid velocity is sub-luminal, leading to $\gamma_{v}\to 1$ and $h\to 1$. Under these approximations, the general relativistic equations of motion reduce to their well-known Newtonian forms
$v\frac{dv}{dr}+\frac{1}{\rho}\frac{dp}{dr}+\frac{GM}{r^{2}}=0,$ (8)
and the mass accretion rate becomes
$\frac{dM}{dt}=4\pi r^{2}\rho v.$ (9)
Most of the studies of accretion onto PBHs have employed these Newtonian equations to model the process within a cosmological context Ali-Haimoud:2016mbv ; Ricotti:2007jk ; Wang:2023qxj ; Tabasi:2022fap . We will show how relativistic corrections in Eqs. (6) and (7) change the outcome compared to the Newtonian one.
We have obtained the general relativistic Hydrodynamic equation in static black hole background. In the following we intend to apply this in the cosmological context where background spacetime is expanding. Exact analysis for such generalization would be to find out the black hole solution in cosmological background which is beyond the scope of our present work. To incorporate the expansion of the universe directly we consider the Friedmann–Lemaître–Robertson–Walker (FLRW) background, parametrized by the scale factor $a(t)$, with the metric $ds^{2}=-dt^{2}+a^{2}(t)(dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta\,d\phi^{2})$, and follow the approximation detailed in Ricotti:2007jk .
@@ II.1 Accretion in the expanding background
where $H~(=\dot{a}/a)$ is the Hubble parameter and $v_{p}~(=a\dot{x})$. The Hubble parameter $H$ satisfies the dynamical equation $H^{2}=\dot{a^{2}}/a^{2}=({1}/{3M_{p}^{2}})\rho$. Where, $\rho$ corresponds to the background energy density of the Universe.
In all practical purpose the BH horizon radius $(r_{s})$ satisfies $r_{s}\ll H^{-1}\rightarrow\epsilon=r_{s}H\ll 1$. Therefore, asymptotic flat black hole approximation which we considered in the earlier discussion can be approximately valid. One can further define a length scale, $r_{*}$ associated with the sphere of influence of a BH within which its gravity tends to dominate over the cosmic expansion. It can be approximately defined as the total background mass contained in the volume of radius $r_{*}$ equals the BH mass, as $M=\frac{4}{3}\pi r_{*}^{3}\rho$. This yields $r_{*}=r_{s}/\epsilon^{2/3}$. Thus, for small $\epsilon$, the gravitational influence extends far beyond the Schwarzschild radius ($r_{*}\gg r_{s}$). Therefore, Within this region, the fluid dynamics are governed by the quasi-static Schwarzschild geometry with Hubble flow being treated as small perturbation to the accretion flow in this limit. However $r>r_{*}$, background can be described by standard cosmological expansion with the scale factor $a(t)$.
Incorporating Eq. (10) into the accretion equations, and integrating Eq. (6) we get,
@@
The integration constant is set by the condition that far away from the BH ($x\rightarrow\infty$), $v\rightarrow 0$, and $\rho\rightarrow\rho_{\infty}\propto a^{-3(1+\omega)}$, which corresponds to the background energy density of the Universe, characterized by the generic isothermal equation of state $p=\omega\rho$. The Hubble parameter $H^{2}=\dot{a^{2}}/a^{2}=({1}/{3M_{p}^{2}})\rho$, can be further expressed as $H=H_{\rm in}\left(a/a_{\rm in}\right)^{-\frac{3}{2}(1+\omega)}$, with index $\rm in$}” being some initial time which we identify later. Here, $M_{p}=1/\sqrt{8\pi G}\simeq 2.435\times 10^{18}\text{Gev}$, being the reduced Planck mass. Utilizing this, the relation between energy density and mass density of the fluid can be obtained by integrating the enthalpy generation Eq. (4) as,
@@
In Fig. 1, we have depicted the various representative solution topologies for the inflowing matter. The physical solutions are the ones where the inflowing matters pass through a critical point and transits smoothly from subsonic to supersonic branch. At the critical point, one needs to satisfy $N=D=0$ condition simultaneously. Solving this, we get the critical point location as $x_{c}=GM(1+3\omega)/(2a\omega)$ and matter velocity at $x_{c}$ as $v^{c}_{p}=\sqrt{\omega}-Hax_{c}$. Inserting these in Eq. (13), the accretion rate at the critical point becomes,
@@
where $\lambda_{c}=(1/4)\omega^{-3/2}(1+3\omega)^{(1+3\omega)/2\omega}$. Following the same procedure for the non-relativistic equations, namely Eq. (8) and Eq. (9), the location of the critical point is determined to be $x_{c}^{\rm NR}=GM/(2a\omega)$. At this point, the peculiar velocity of the matter is given by $v_{p,c}^{\rm NR}=\sqrt{\omega}-GMH/(2\omega)$. These values, in turn, define the accretion rate expressed in Eq. (15), with the non-relativistic dimensionless critical eigenvalue $\lambda_{c}^{\rm NR}$ taks the form $\lambda_{c}^{\rm NR}=(1/4)\exp(3/2)\omega^{-3/2}$, which is consistent with the results presented in Ali-Haimoud:2016mbv for an isothermal equation of state. Integrating Eq. (15), we obtain
$\frac{M}{M_{\mathrm{in}}}=\left[1-\frac{\lambda_{c}\gamma}{2(1+\omega)}\left(\frac{\frac{3}{2}H_{\rm in}(1+\omega)(t-t_{\rm in})}{\frac{3}{2}H_{\rm in}(1+\omega)(t-t_{\rm in})+1}\right)\right]^{-1}\;$ (16)
where $M_{\mathrm{in}}~(=4\pi\gamma M_{p}^{2}H_{\rm in}^{-1})$ is the initial mass of a PBH with $H_{\rm in}$ being the Hubble parameter at the formation time $t=t_{\rm in}$. Here, $\gamma~(=\omega^{3/2})$ is the collapse fraction to form PBH Carr & Hawking (1974). From Eq. (16), we indeed see that PBH gains mass due to accretion and quickly saturates into (see Fig. 2),
@@ III Hawking evaporation
For example, assuming a PBH with a mass of 10 g formed in a radiation-dominated background ($\omega=1/3$), it completely evaporates at approximately $\tau_{\rm BH}^{\rm ev}\simeq 8.8\times 10^{17}M_{p}^{-1}$. This highlights the wide separation between the timescales of the two different physical processes of our interest, namely accretion and evaporation ($\tau_{\rm BH}^{\rm ev}\gg\tau_{\rm BH}^{\rm acc}$), as depicted in Fig. 2.
@@
[…]
@@ IV PBH mass evolution
The evolution of PBH mass, accounting for both relativistic (blue) and non-relativistic (red) accretion in conjunction with Hawking evaporation, is illustrated in Fig. 2. Results are presented for two initial masses, namely $10\,\mathrm{g}$ (dashed lines) and $10^{8}\,\mathrm{g}$ (solid lines), respectively.
@@ V BBN Constraint
PBHs evaporating during the era of Big Bang Nucleosynthesis (BBN) can inject additional relativistic degrees of freedom, potentially altering the primordial element abundances predicted by BBN. To preserve the observed consistency of BBN results, PBHs must evaporate before the onset of BBN. This requirement immediately suggests that the PBH mass within the range of $10^{8}\,\text{g}\lesssim M_{\rm BH}\lesssim 10^{13}\,\text{g}$ Carr:2020gox ; Keith:2020jww ; Carr:2009jm should not form in the early universe. However, when PBH accretion is taken into account, these mass constraints are naturally altered.
Figure 2: PBH mass ($M$) evolution as a function of scale factor ($a$) for initial masses of 1010 g (dashed) and $10^{8}$ g (solid), comparing relativistic (green) and non-relativistic (red) accretion along with Hawking evaporation.
@@
$T_{\rm ev}\simeq\sqrt{\frac{3\epsilon}{2}}\left(\frac{\pi^{2}g_{\ast}}{30}\right)^{-\frac{1}{4}}\left(\frac{M_{p}}{M_{\mathrm{in}}}\right)^{\frac{3}{2}}\left(1-\frac{3\gamma\lambda_{c}}{8}\right)^{\frac{3}{2}}M_{p},$~~ (27)
@@
In the above expression, we consider $g_{*}=106.75$ at the time of PBH formation. Equation (28) clearly demonstrates the impact of accretion on the BBN limit for the minimum PBH mass. For example, if we set $\lambda_{c}=0$ (without accretion), we recover the well known BBN bound on PBH initial mass $M_{\rm in}\simeq 3.94\times 10^{8}\,\mathrm{g}$. On contrary, relativistic accretion ($\lambda_{c}\simeq 10.4$) clearly shifts such bound to one order lower value, $M_{\rm in}\simeq 8.6\times 10^{7}\,\mathrm{g}$. Our analytic estimates match well with the numerical ones illustrated in Fig. 3, and further depicts the dependence of the initial PBH mass on $T_{\rm ev}$ for different accretion scenarios.
Figure 3: Initial PBH mass ($M_{\rm in}$) as a function of evaporation temperature ($T_{\rm ev}$), shown for three cases namely, blue, red, and green curves representing without, with non-relativistic, and with relativistic accretion respectively.